OpenStudy (anonymous):

Steps to factorize a^4+4b^4

OpenStudy (anonymous):

you can take a square like... \[(a^{2})^{2}+(2b ^{2})^{2}\]

OpenStudy (anonymous):

you have to complete the square

OpenStudy (anonymous):

1. what are the factors that will get you the answers? a^2 * a^2 and 2b^2 and 2b^2 2. Now, since there is no middle term and the terms are both positive, you know that it needs to be + * -. Ergo 3. (a^2 + 2b^2) (a^2 - 2b^2)

OpenStudy (anonymous):

This pattern is also called the sum of two squares

OpenStudy (mr.math):

\[a^4+4b^4=(a^2)^2+(2b)^2+4a^2b^2-4a^2b^2=(a^2+2b^2)^2-4a^2b^2\] \[=(a^2+2b^2)-(2ab)^2.\] Use the difference between squares to factorize the above expression and you're done.

OpenStudy (mr.math):

Sorry typo, second line should be \[=(a^2+2b^2)^2-(2ab)^2.\]

OpenStudy (mr.math):

@dreamergurl your answer would be correct if the expression was \(a^4-4b^4\).

OpenStudy (anonymous):

ugh you're right. Thanks, Mr. Math.

OpenStudy (anonymous):

Wait, how did you get my name to turn blue?

OpenStudy (mr.math):

I just added "@" before it as you can see :-)

OpenStudy (mr.math):

Nice picture and nice name by the way.